OpenStudy (anonymous):

Answer 88 m2 136 m2 100 m2 70 m2

OpenStudy (anonymous):

all of the above

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

none of them

OpenStudy (angela210793):

O.o

OpenStudy (anonymous):

sorry forgot to put the picture up

OpenStudy (angela210793):

u want the area or wht?

OpenStudy (anonymous):

the area

OpenStudy (anonymous):

88 square meter

OpenStudy (angela210793):

A=(B+b)*h/2 + b*h=(20+12)*3/2+20*4/2=48+40=88m^2

OpenStudy (angela210793):

(B+b)*h/2=Area of trapezoid and b*h/2 area of triangle